# BEN: Using Confidence-Guided Matting for Dichotomous Image Segmentation

### v1

@@ Abstract
Current approaches to dichotomous image segmentation (DIS) treat image matting and object segmentation as fundamentally different tasks. As improvements in image segmentation become increasingly challenging to achieve, combining image matting and grayscale segmentation techniques offers promising new directions for architectural innovation. Inspired by the possibility of aligning these two model tasks, we propose a new architectural approach for DIS called Confidence-Guided Matting (CGM). We created the first CGM model called Background Erase Network (BEN). BEN is comprised of two components: BEN Base for initial segmentation and BEN Refiner for confidence refinement. Our approach achieves substantial improvements over current state-of-the-art methods on the DIS5K validation dataset, demonstrating that matting-based refinement can significantly enhance segmentation quality. This work opens new possibilities for cross-pollination between matting and segmentation techniques in computer vision.
@@ 1 Introduction
Image segmentation is used in a wide range of tasks, including autonomous vehicles [1, 2], and medical imaging [3, 4, 5]. Dichotomous image segmentation (DIS) attempts to correctly segment foreground objects from real images. The DIS5K [6] dataset contains 5,470 high-resolution images and their corresponding pixel-accurate masks. This dataset is comprised of a wide range of diverse categories, from musical instruments to automobiles. The DIS5K has become a vital benchmark for understanding segmentation performance in real-world scenarios.
Recent efforts have lead to significant improvements on the DIS5K through novel architectures. The Multi-view Aggregation Network (MVANet) [7] achieved state-of-the-art results using multi-view learning and a Swin backbone  [8]. However, transformer-based approaches face limitations due to the computational cost of attention in Vision Transformer (ViT) [9] and Swin backbone architectures. Additionally, these limitations make batch edge prediction impractical. When encountering data distributions significantly different from their training set, these ViT and Swin backbone models often display reduced confidence in foreground-background classification, leading to unwanted matting effects caused by uncertainty. The DiffDIS [10] demonstrated that incorporating edge prediction alongside foreground segmentation improves the accuracy of its final output. DiffDIS employs a Variational Auto Encoder (VAE) [11] in order to enable the diffusion model to operate on high-resolution images in a compressed latent space before they are reconstructed to their input resolution. This approach supports scalable batch sizes, facilitating precise edge and foreground prediction.
We propose Confidence-Guided Matting (CGM), an architectural approach that bridges the gap between segmentation and matting to enable coherent foreground segmentation. Current matting approaches generate trimaps manually. These trimaps are used to predict fine-grained details, focusing on regions where the foreground meets the background. Trimaps consist of three regions: black (background), white (foreground), and gray (unknown). The gray area of the trimap is updated with the foreground or background prediction by a matting model.
Our method innovates on existing matting techniques by dynamically generating trimaps based on the confidence levels from the base model’s predictions. These confidence trimaps allow a refiner network to focus on uncertain areas, where the foreground transitions into the background or regions with complex, fine-grained details. Our Background Erase Network (BEN) incorporates this approach by employing a base network for initial segmentation and a refiner network for confidence-guided refinement. Together, these components work in concert, as depicted in Figure 1, to achieve new state-of-the-art accuracy on the DIS5K validation dataset.
Figure 1: CGM Pipeline: Shows the process flow from input image through Base Model, Confidence Trimap Algorithm, and Refiner Model stages to produce the final segmentation mask.
@@ 2.1 Advances in Dichotomous Image Segmentation
The Multi-view Aggregation Network (MVANet) [7] has shown significant promise in DIS using a multi-view learning approach. The MVANet separates the images into global and local patches. These patches are concatenated batch-wise and given to a Swin Transformer  [8] backbone. The Swin Transformer generates different hierarchical representations of each of the patches. The patches share information with each other using Multi-view Complementary Refinement and Multi-view Complementary Localization blocks and are up-sampled in the process. This process ultimately aggregates the information into one final output. This architectural approach shows state-of-the-art performance while keeping a modest 94 million parameters. In this paper, we take the core ideas of the MVANet architecture and apply them to our BEN Base model.
Diffusion DIS (DiffDIS) [10] attempts to shift to a diffusion-based approach, as diffusion has played a crucial role in computer vision  [12]. The DiffDIS allows for a multi-batch estimation of the segmentation and edge prediction. Edge prediction alongside segmentation significantly increases the detail of the foreground’s edges. Along with other innovations, this DiffDIS model surpasses the MVANet model on the DIS5K validation dataset. The key contribution from the DiffDIS to this paper is the insight that fine-grained edge accuracy will increase when predicting the edges of any given foreground scene. This idea is one of our key motivations for CGM approach.
@@ 2.2 Deep Learning Image Matting
Image matting leverages alpha mattes to discern the foreground from the background. The ambiguous areas often describe complicated scenes like hair or smoke. There has been a shift from convolution neural network approaches to transformer-based architectures. Vision Transformer-based matting models have shown state-of-the-art results by incorporating attention mechanisms [13]. The models show great efficacy in using trimaps to guide foreground refinement. These techniques pave the way for confidence-based trimap generation, hence our work unifying them with CGM on the DIS5K.
@@ 3.1 BEN Base
The BEN Base architecture closely resembles that of the MVANet but with notable changes. We changed the activation function and normalization when redesigning the MVANet for our base model. The MVANet was trained with a batch size of one. With only a batch size of one, we opted to use instance normalization instead of batch normalization, as batch normalization is unstable for training with a small batch size [14]. We also wanted to leverage the power of the Gaussian Error Linear Unit [15] as our activation function rather than the Rectified Linear Unit and Parametric Rectified Linear Unit used in the original MVANet.
@@ 3.2 Loss Function
We use three metrics for BEN Base’s loss function: Weighted BCE, Weighted IoU, and weighted structural similarity (SSIM). We define a Structure Loss to measure the difference between a predicted segmentation (logits) and its ground-truth mask. Our loss closely resembles that of the BiRefNet [16]. This loss leverages all three of the previous equations. Formally written as:
$\text{StructureLoss}(p,y)\;=\;w_{\text{BCE}}\cdot\text{WBCE}(p,y)\;+\;w_{\text{IOU}}\cdot\text{WIOU}(\sigma(p),y)\;+\;w_{\text{SSIM}}\cdot\text{SSIM}(\sigma(p),y),$ (1)
@@
$\mathcal{L}_{\text{local}}=\sum_{i=1}^{6}\text{StructureLoss}(s_{i},y_{i})$ (2)
@@
$\mathcal{L}_{\text{token}}=\sum_{i=1}^{4}\text{StructureLoss}(t_{i},y_{i}),$ (4)
where $s_{i}$ are side outputs and includes the the final output, $g_{i}$ are global outputs. where $t_{i}$ are token outputs at different scales. Building on these scale-specific terms, we define a combined loss that balances local, global, and token-level predictions. This normal combined loss is then defined by:
$\mathcal{L}_{\text{combined}}=0.3\,\mathcal{L}_{\text{global}}\;+\;0.3\,\mathcal{L}_{\text{token}}\;+\;\mathcal{L}_{\text{local}}.$ (5)
3.3 Confidence-Guided Matting
To address the limitations of matting’s current application in DIS techniques, we propose Confidence-Guided Matting (CGM). CGM allows a base prediction to be updated depending on the model’s confidence that a given pixel is a foreground or background pixel. Unlike static trimaps that are created by hand labeling or set thresholds, our trimaps are generated purely by the base model’s confidence that a pixel is a part of the foreground or background. The refiner network subsequently learns how to optimize the uncertain regions leading to improved segmentation accuracy.
To amalgamate the base and refiner model, we generate confidence trimaps from the base model’s sigmoid prediction. These confidence trimaps are then passed to the refiner for the final prediction. To simplify the explanation of the algorithm, we’ve formally written two stages: initial trimap generation and iterative refinement. The first stage involves generating an initial trimap from the base model’s sigmoid predictions. The second stage refines the initial trimap using iterative updates to ensure adequate uncertainty regions. To see the Python implementation, see Appendix A.
Input : Predicted mask $P$, dimensions $H=W=1024$
@@ 3.3 Confidence-Guided Matting
1 Initialize:
2 $t_{h}\leftarrow 0.90$, $t_{l}\leftarrow 0.10$, $M\leftarrow\sigma(P)$
3
4foreach $(i,j)\in H\times W$ do
@@ 14             $T_{i,j}\leftarrow 128$
Algorithm 1 Initial Trimap Generation.
Input : Initial Trimap $T$, sigmoid mask $M$, thresholds $t_{l}$, $t_{h}$, and constants $t_{\text{min}}$, $t_{\text{max}}$, $s$, $N_{\text{min}}$
Output : Refined Trimap $T\in\{0,128,255\}^{H\times W}$
1 $n_{g}\leftarrow\big{|}\{(i,j):T_{i,j}=128\}\big{|}$
2
3while $n_{g}<N_{\text{min}}$ and $(t_{l}>t_{\text{min}}$ or $t_{h}<t_{\text{max}})$ do
4       $t_{l}\leftarrow\max(t_{l}-s,t_{\text{min}})$
5       $t_{h}\leftarrow\min(t_{h}+s,t_{\text{max}})$
6
7      foreach $(i,j)\in H\times W$ do
8             if $M_{i,j}\geq t_{h}$ then
9                   $T_{i,j}\leftarrow 255$
10
11             end if
12            else if $M_{i,j}\leq t_{l}$ then
13                   $T_{i,j}\leftarrow 0$
14
15             end if
16            else
17                   $T_{i,j}\leftarrow 128$
18
19             end if
20
21       end foreach
22      $n_{g}\leftarrow\big{|}\{(i,j):T_{i,j}=128\}\big{|}$
23
24 end while
return $T$
Algorithm 2 Trimap Refinement.
Table 1: Quantitative comparison of DIS5K with common representative methods to compare our models to existing DIS models. $\uparrow$ represents “higher is better,” and $\downarrow$ represents “lower is better.” The best score is in bold. External scores are from  [10].
@@ 17                   $T_{i,j}\leftarrow 128$
$F_{\beta}^{max}\uparrow$ 0.889 0.897 0.913 0.844 0.918 0.9234 0.9188
$F_{\beta}^{\omega}\uparrow$ 0.834 0.863 0.856 0.824 0.888 0.8708 0.8956
$E_{\phi}^{m}\uparrow$ 0.914 0.937 0.938 0.924 0.948 0.9346 0.9584
$S_{m}\uparrow$ 0.900 0.905 0.905 0.863 0.904 0.9161 0.9166
$\mathcal{M}\downarrow$ 0.042 0.036 0.036 0.044 0.029 0.0309 0.0270
Table 2: Comprehensive comparison of our BEN models with MVANet. $\uparrow$ indicates “higher is better” and $\downarrow$ indicates “lower is better.” The best score is in bold.
@@
$\mathrm{MAE}\downarrow$ 0.0356 0.0309 0.0270
$\mathrm{Dice}\uparrow$ 0.8691 0.8806 0.8989
$\mathrm{IoU}\uparrow$ 0.8069 0.8371 0.8506
$\mathrm{BER}\downarrow$ 0.0647 0.0516 0.0496
$\mathrm{Acc}\uparrow$ 0.9656 0.9718 0.9740
4 Experiments
4.1 Data and Evaluation Metrics
To leverage both commercial interest and the open-source community, we train on the DIS5K dataset and preserve the DIS5K validation set for fair comparison with other publicly available models. We evaluate our results using widely adopted metrics as follows: Max F-measure ($F_{\max}^{\beta}$[17], Weighted F-measure ($F_{\omega}^{\beta}$[18], Structural Similarity Measure ($S_{m}$[19], E-measure ($E_{m}^{\phi}$[20], Mean Absolute Error (MAE, $M$[17], Dice Coefficient (Dice) [21], Intersection-over-Union (IoU) [22], Balanced Error Rate (BER) [23], and Accuracy (Acc).
4.2 Implementation Details
For our confidence trimap generation we set the minimum number of gray pixels to 60,000 for a 1024 by 1024 base prediction. We set our initial high and low confidence values at 0.9 and 0.1. Their max and min values are set to 0.97 and 0.03. The confidence values are also incremented and decremented by a step size of 0.001. It should be noted that the BEN Base is first trained to reach its highest accuracy possible and the Refiner model is subsequently trained on the best version of the BEN Base.
@@ 4.3 Qualitative Results
To evaluate our model, we examined the data from 5 different models, InSPyReNet [24], BiRefNet [16], MVANet [7], GenPercept [25] and DiffDIS [10] Table 1. The BEN family occupy the top slot for each of the DIS 5k measurement methods but BEN Base on the ($F_{\max}^{\beta}$) performs better than than BEN Base+Refiner. To do a deeper examination of our performance we expanded our model evaluation with more metrics and ran the MVANet locally to act as a control depicted in Table 2. BEN Base + Refiner shows impressive dominance over the MVANet and BEN Base as well.
Input GT Confidence
Trimap
BEN Base
+ Refiner
BEN Base MVANet
Figure 2: Qualitative results from DIS5K validation
4.4 Quantitative Results
To intuitively understand how the refiner is updating the base predictions we selected outputs from the confidence trimap algorithm, BEN Base, and BEN Base + Refiner that are in the training dataset. As shown in Figure 2 the refiner is correctly able to remove and refine edges of the base models prediction. This combination significantly outperforms the MVANet as it is unable to correctly define foreground objects and attend to fine details.
@@ 5 Conclusion
In this paper, we examine the union of image foreground segmentation and image matting with the creation of BEN. We leverage CGM to allow the Refiner sufficient control over uncertain areas from BEN Base’s initial prediction. This powerful combination shows significant improvements over existing segmentation methods on the DIS5K validation dataset. Our work supports further exploration of confidence-based matting techniques in image segmentation.
Appendix A PyTorch Implementation of Confidence Trimap Generation
Basic PyTorch implementation of combined Algorithm 1 and Algorithm 2 is shown in Listing 1. This implementation provides an efficient vectorized version of the confidence trimap generation algorithm.
Listing 1: PyTorch Implementation of confidence trimap generation
1def generate_trimap(pred):
2 """
3 Generate a confidence-based trimap from a predicted mask.
4 Args:
5 pred: Predicted mask tensor
6 min_pixels: Minimum number of gray pixels required
7 Returns:
8 trimap: Generated trimap tensor
9 """
10 # Initialize thresholds
11
12 min_pixels=60000
13 t_high = 0.90
14 t_low = 0.10
15 t_min = 0.03 # Lower bound
16 t_max = 0.97 # Upper bound
17 step = 0.001 # Adjustment size
18 mask = pred.sigmoid() # Apply sigmoid to prediction
19
20 # Generate initial trimap
21 trimap = torch.where(
22 mask >= t_high,
23 mask.new_tensor(255.0),
24 torch.where(
25 mask <= t_low,
26 mask.new_tensor(0.0),
27 mask.new_tensor(128.0)
28 )
29 )
30
31 # Count gray pixels and adjust thresholds
32 n_gray = (trimap == 128).sum().item()
33 while n_gray < min_pixels:
34 t_low = max(t_low - step, t_min)
35 t_high = min(t_high + step, t_max)
36
37 if (t_low <= t_min and t_high >= t_max):
38 break # Exit if bounds reached
39
40 trimap = torch.where(
41 mask >= t_high,
42 trimap.new_tensor(255.0),
43 torch.where(
44 mask <= t_low,
45 trimap.new_tensor(0.0),
46 trimap.new_tensor(128.0)
47 )
48 )
49 n_gray = (trimap == 128).sum().item()
50
51 return trimap

### v2

@@ Abstract
Current approaches to dichotomous image segmentation (DIS) treat image matting and object segmentation as fundamentally different tasks. As improvements in image segmentation become increasingly challenging to achieve, combining image matting and grayscale segmentation techniques offers promising new directions for architectural innovation. Inspired by the possibility of aligning these two model tasks, we propose a new architectural approach for DIS called Confidence-Guided Matting (CGM). We created the first CGM model called Background Erase Network (BEN). BEN consists of two components: BEN Base for initial segmentation and BEN Refiner for confidence-based refinement. Our approach achieves substantial improvements over current state-of-the-art methods on the DIS5K validation dataset, demonstrating that matting-based refinement can significantly enhance segmentation quality. This work introduces a new paradigm for integrating matting and segmentation techniques, improving fine-grained object boundary prediction in computer vision.
Figure 1: Shows the process flow from input image through Base Model, Confidence Trimap Algorithm, and Refiner Model stages to produce the final segmentation mask.
@@ 1 Introduction
Image segmentation is used in a wide range of tasks, including autonomous vehicle navigation [2, 1] and medical imaging [19, 24, 3]. Dichotomous image segmentation (DIS) is the segmentation of foreground objects from real images. The DIS5K [14] dataset contains 5,470 high-resolution images and their corresponding pixel-accurate masks. This dataset is comprised of a wide range of diverse categories, from musical instruments to automobiles. The DIS5K dataset has become a vital benchmark for understanding segmentation performance in real-world scenarios.
Recent efforts have led to significant accuracy improvements on the DIS5K dataset through novel architectures. The Multi-view Aggregation Network (MVANet) [23] achieved state-of-the-art results using multi-view learning and a Swin backbone [11]. However, transformer-based approaches face limitations due to the computational cost of attention in Vision Transformer (ViT) [4] and Swin backbone architectures. Additionally, these limitations make batch edge prediction impractical. When encountering data distributions significantly different from their training set, these ViT and Swin backbone models often display reduced confidence in foreground-background classification, leading to unwanted matting effects. The DiffDIS [22] demonstrated that incorporating edge prediction alongside foreground segmentation improves the accuracy of the final output. DiffDIS employs a Variational Auto Encoder (VAE) [9] in order to enable the diffusion model to operate on high-resolution images in a compressed latent space before they are reconstructed to their input resolution. This approach supports scalable batch sizes, facilitating precise edge and foreground prediction.
We propose Confidence-Guided Matting (CGM), an architectural approach that bridges the gap between segmentation and matting, enabling improved segmentation performance. Current matting approaches generate trimaps manually. These trimaps are used to predict fine-grained details, focusing on regions where the foreground meets the background. Trimaps consist of three regions: black (background), white (foreground), and gray (unknown). The gray area of the trimap is updated with the foreground or background prediction by a matting model.
Our method innovates on existing matting techniques by dynamically generating trimaps based on the confidence levels from the base model’s predictions. These confidence trimaps allow a refiner network to focus on uncertain areas, where the foreground transitions into the background or regions with complex, fine-grained details. Our Background Erase Network (BEN) incorporates this approach by employing a base network for initial segmentation and a refiner network for confidence-guided refinement. Together, these components work in concert, as depicted in Figure 1, to achieve new state-of-the-art accuracy on the DIS5K validation dataset (DIS-VD).
@@ 2.1 Advances in Dichotomous Image Segmentation
The Multi-view Aggregation Network (MVANet) [23] has shown significant promise in DIS by utilizing a multi-view learning approach. The MVANet separates the images into global and local patches, which are concatenated batch-wise and then inputted to a Swin Transformer [11] backbone. The Swin Transformer generates hierarchical representations of each of the patches, which then share information with one another using Multi-view Complementary Refinement and Multi-view Complementary Localization blocks, which are up-sampled in the process. This process ultimately aggregates the information into one final output. This architectural approach shows state-of-the-art performance at a modest 94 million parameters. We expand the core ideas of the MVANet architecture to create our BEN Base model.
Diffusion DIS (DiffDIS) [22] attempts to shift to a diffusion-based approach, as diffusion has played a crucial role in computer vision [16]. The DiffDIS allows for a multi-batch estimation of the segmentation and edge prediction. Edge prediction alongside segmentation significantly increases the detail of the foreground’s edges. Along with other innovations, this DiffDIS model surpasses the MVANet model on the DIS5K validation dataset. The key contribution from the DiffDIS to our approach is the insight that fine-grained edge accuracy will increase when predicting the edges of any given foreground scene. This idea is one of our key motivations for CGM approach.
@@ 2.2 Deep Learning Image Matting
Image matting leverages alpha mattes to discern the foreground from the background. The ambiguous areas are often complicated scenes such as hair or smoke. There has been a shift from convolution neural network approaches to transformer-based architectures. Vision Transformer-based matting models have shown state-of-the-art results by incorporating attention mechanisms [21]. The models show great efficacy in using trimaps to guide foreground refinement. These techniques pave the way for confidence-based trimap generation, hence our work unifying them with CGM on the DIS5K dataset.
@@ 3.1 BEN Base
The BEN Base architecture resembles that of the MVANet but with notable changes. We modified the activation function and normalization when redesigning the MVANet for our base model. MVANet was originally trained with a batch size of one; therefore, we use instance normalization instead of batch normalization in BEN Base, as batch normalization is known to be unstable for small batch sizes [18]. We adopt Gaussian Error Linear Unit [7] as the activation function rather than the Rectified Linear Unit and Parametric Rectified Linear Unit used in the original MVANet.
@@ 3.2 Loss Function
We use three metrics for BEN Base’s loss function: weighted BCE (WBCE), weighted IoU (WIOU), and weighted structural similarity (SSIM). We define a Structure Loss to measure the difference between a predicted segmentation (logits) and its ground-truth mask. Our loss closely resembles that of the BiRefNet [25]. This loss leverages all three of the previous equations. Formally written as:
$\begin{split}\text{StructureLoss}(p,y)&=w_{\text{BCE}}\cdot\text{WBCE}(p,y)\\ &\quad+w_{\text{IOU}}\cdot\text{WIOU}(\sigma(p),y)\\ &\quad+w_{\text{SSIM}}\cdot\text{SSIM}(\sigma(p),y),\end{split}$ (1)
@@
$\mathcal{L}_{\text{local}}$ $=\sum_{i=1}^{6}\text{StructureLoss}(s_{i},y_{i}),$ (2)
@@
$\mathcal{L}_{\text{token}}$ $=\sum_{i=1}^{4}\text{StructureLoss}(t_{i},y_{i})$ (4)
where $s_{i}$ are side outputs and includes the final output, $g_{i}$ are global outputs. where $t_{i}$ are token outputs at different scales. Building on these scale-specific terms, we define a combined loss that balances local, global, and token-level predictions. This normal combined loss is then defined by:
$\mathcal{L}_{\text{combined}}=0.3\,\mathcal{L}_{\text{global}}\;+\;0.3\,\mathcal{L}_{\text{token}}\;+\;\mathcal{L}_{\text{local}}$ (5)
Input : Predicted mask $P$, $H=W=1024$
@@
1 $t_{h}\leftarrow 0.95,\;t_{l}\leftarrow 0.05,\;M\leftarrow\sigma(P)$
2
3foreach $(i,j)\in H\times W$ do
@@ 11    $T_{i,j}\leftarrow 128$
Algorithm 1 Trimap Generation
3.3 Confidence-Guided Matting
To address the limitations of matting’s current application in DIS techniques, we propose Confidence-Guided Matting (CGM). CGM allows a base prediction to be updated depending on the model’s confidence that a given pixel is a foreground or background pixel. Unlike traditional trimaps, which require manual labeling or predefined thresholds, our trimaps are dynamically generated based on the base model’s confidence in each pixel’s classification. The refiner network subsequently learns how to optimize the uncertain regions, which leads to improved segmentation accuracy. To integrate the base and refiner model, we generate confidence trimaps from the base model’s sigmoid prediction. These confidence trimaps are then passed to the refiner for the final prediction. The trimap generation process is formally presented in Algorithm 1.
Input GT Confidence Trimap BEN Base+ Refiner BEN Base MVANet
Figure 2: Qualitative results from DIS5K validation dataset
@@ 3.3 Confidence-Guided Matting
$F_{\beta}^{max}\uparrow$ 0.889 0.897 0.913 0.844 0.918 0.923 0.919
$F_{\beta}^{\omega}\uparrow$ 0.834 0.863 0.856 0.824 0.888 0.871 0.896
$E_{\phi}^{m}\uparrow$ 0.914 0.937 0.938 0.924 0.948 0.935 0.959
$S_{m}\uparrow$ 0.900 0.905 0.905 0.863 0.904 0.916 0.917
$\mathcal{M}\downarrow$ 0.042 0.036 0.036 0.044 0.029 0.031 0.027
Table 1: Quantitative comparison of DIS5K validation dataset with common representative methods to compare our models to existing DIS models. $\uparrow$ represents "higher is better," and $\downarrow$ represents "lower is better". The best score is indicated in bold. External scores are from [22].
4 Experiments
4.1 Data and Evaluation Metrics
To leverage both commercial interest and the open-source community, we train on the DIS5K dataset and preserve the DIS5K validation set for fair comparison with other publicly available models. We evaluate our results using widely adopted metrics as follows: Max F-measure ($F_{\beta}^{max}$) [13], Weighted F-measure ($F_{\beta}^{\omega}$) [12], E-measure ($E_{\phi}^{m}$) [6], Structural Similarity Measure ($S_{m}$) [5], Mean Absolute Error (MAE, $\mathcal{M}$) [13], Dice Coefficient (Dice) [17], Intersection over Union (IoU) [15], Balanced Error Rate (BER) [10], and Accuracy (Acc).
@@ 4.1 Data and Evaluation Metrics
[…]
Table 2: Comprehensive comparison of our BEN models with MVANet.
4.2 Quantitative Results
To evaluate our model, we examined the data from 5 different models, InSPyReNet [8], BiRefNet [25], MVANet [23], GenPercept [20] and DiffDIS [22] Table 1. The BEN family occupies the top slot for each of the DIS 5k measurement methods. Based on the ($F_{\beta}^{max}$) metric, BEN Base alone slightly outperforms BEN Base+Refiner. To perform a deeper examination of our performance, we expanded our evaluation with additional metrics and used MVANet as a control depicted in Table 2. BEN Base + Refiner significantly outperforms MVANet and BEN Base as well.
@@ 4.3 Qualitative Results
To intuitively understand how the refiner is updating the base predictions we selected outputs from the confidence trimap algorithm, BEN Base, and BEN Base + Refiner that are in the training dataset. As shown in Figure 2, the Refiner is correctly able to remove and refine edges of the base models prediction. This combination significantly outperforms the MVANet as it is unable to correctly define foreground objects and attend to fine details.
4.4 Ablation Test
DIS-VD
Low High MAE Dice IoU BER Acc
0.45 0.55 0.029 0.895 0.842 0.049 0.972
0.35 0.65 0.028 0.896 0.845 0.050 0.973
0.25 0.75 0.028 0.897 0.847 0.050 0.973
0.15 0.85 0.027 0.898 0.849 0.050 0.974
0.05 0.95 0.027 0.899 0.851 0.049 0.974
0.01 0.99 0.027 0.899 0.851 0.050 0.974
0.005 0.995 0.027 0.897 0.848 0.051 0.974
Table 3: Comprehensive comparison of various refinement settings for BEN_Base+Refiner on the DIS-VD. $\uparrow$ indicates "higher is better" and $\downarrow$ indicates "lower is better". The best scores are indicated in bold.
We examine the impact of modifying the low and high confidence thresholds in the trimap generation. In Table 3, the comparison of the thresholds shows the efficacy of the proposed 0.05 and 0.95 thresholds. These thresholds show the best trade-off between the refiner’s control of the output versus the base prediction’s control.
@@ 5 Conclusion
In this paper, we examine the union of image foreground segmentation and image matting with the creation of BEN. We leverage CGM to dynamically generate confidence trimaps that allow the Refiner sufficient control over uncertain areas from BEN Base’s initial prediction. Automating the trimap generation enables the Refiner to learn uncertain regions. This powerful combination shows significant improvements over existing segmentation methods on the DIS5K validation dataset. This introduces a novel approach to segmentation, unifying matting with high-precision segmentation. We plan to extend this work to other image segmentation tasks in future work. Our work supports further exploration of confidence-based matting techniques in precise image segmentation.